OpenStudy (anonymous):

Solve the equation. 3|2x - 5| = 12

OpenStudy (bahrom7893):

divide both sides by 3: |2x-5|=4

OpenStudy (bahrom7893):

since absolute value makes anything positive: (2x-5)=4 and (2x-5)=-4

OpenStudy (bahrom7893):

solve for x in both cases.

OpenStudy (anonymous):

thanks . .